OpenStudy (anonymous):

what is the sum of the first five terms of a geometric series with a1=20 and r=1/4

OpenStudy (anonymous):

So...it goes like this: The sum of the first n terms of a geometric sequence is given by sn = a1 + a2 + a3 + ... an = a1 (1 - rn) / (1 - r)The sum S of an infinite (n approaches infinity) geometric sequence and when |r| <1 is given byS = a1 / (1 - r) .............SOLVE.........

OpenStudy (anonymous):

That doesn't help me. But thanks anyway.

OpenStudy (anonymous):

this formula should be part of your course Σ = {20 [1 - (1/4)^5]} / (1 - 1/4) Does this help? :O

OpenStudy (anonymous):

No. I just want the answer. LOL

OpenStudy (anonymous):

ok a sec...

OpenStudy (anonymous):

is the answer 26.56?

OpenStudy (anonymous):

no. it was 1705/64

OpenStudy (anonymous):

yes same comes when you simplify....just a slight change 25.64

OpenStudy (anonymous):

Sure.but where is my medal?LOL :D

OpenStudy (anonymous):

If your starting salary were $50,000 and you received a 4% increase at the end of every year for 15 years, what would be the total amount, in dollars, you would have earned over the first 16 years that you worked?